A quantity of gas has a volume of 15 liters at 120.5 degrees Celsius. What is its volume at 120.5 K?
1
Expert's answer
2016-04-25T12:16:05-0400
V/T = V1/ T1
V1 = V*T1/T = 15 * 120.5 / (120.5+273.15) = 4.6 (l)

Answer
V1 = 4.6 l